OpenStudy (anonymous):

ILL LOVE YOU FOREVER IF YOU ANSWER THIS In the early 1970's, Canada Post started using six-character postal codes. Each postal code uses three letters and three digits in an alternating pattern. How many different Canadian postal codes are possible?

OpenStudy (anonymous):

AND OR THIS ONE :))

OpenStudy (here_to_help15):

This 1 is a permutation problem. The big difference between permutation and combination is that permutation cares about the order of things while combination does not. We need to alternate letters and numbers. So there's two initial cases. The first one is when we start with a letter first. There are 26 * 10 * 26 * 10 * 26 * 10 = 17,576,000 different postal codes if we started with a letter first. If we started with a number first then there are 10 * 26 * 10 * 26 * 10 * 26 = 17,576,000 different postal codes. Therefore, the total number of postal codes are 17,576,000 + 17,576,000 = 35,152,000.

OpenStudy (mathstudent55):

I answered this question yesterday.

OpenStudy (anonymous):

I think I didn't save it haha sorry @mathstudent55

OpenStudy (mathstudent55):